OpenStudy (anonymous):

Write the series in summation notation: -3 + 6 - 9 + 12 - 15

OpenStudy (anonymous):

-3+6-9+12-15 \[=\left( -1 \right)^1\left( 3*1 \right)+\left( -1 \right)^2\left( 3*2 \right)+...+\left( -1 \right)^5\left( 3*5 \right)\] \[=\sum_{1}^{5}\left( -1 \right)^n \left( 3n \right)\]